# A Note on Output Length of One-Way State Generators

### v1

@@ Abstract
We study output length of one-way state generators (OWSGs) and their weaker variants.
@@
    Let $\epsilon$-OWSGs be a parameterized variant of OWSGs where a quantum polynomial-time adversary’s advantage is at most $\epsilon$. For any constant $\epsilon>0$, we construct $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit outputs assuming the existence of subexponentially secure OWFs. We show that this is almost tight by proving that there do not exist $O(1)$-OWSGs with $(\log\log\lambda)/2+O(1)$-qubit outputs.
@@
    We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs. We construct weak OWSGs with $m$-qubit outputs for any $m=\omega(1)$ assuming the existence of exponentially secure injective OWFs with linear expansion. We show that this is tight by proving that there do not exist weak OWSGs with $O(1)$-qubit outputs.
@@ 1 Introduction
In classical cryptography, the existence of one-way functions (OWFs) is the minimum assumption, because many primitives (such as secret-key encryption (SKE), commitments, zero-knowledge, pseudorandom generators (PRGs), pseudorandom functions (PRFs), and digital signatures) are equivalent to OWFs, and almost all cryptographic primitives imply OWFs [LR86, IL89, ILL89]. On the other hand, in quantum cryptography where quantum computing and quantum communications are available, recent studies suggest that OWFs would not be the minimum assumption, and many new replacements of OWFs have been introduced. For example, pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22b, MY22a], and EFI [BCQ23] are quantum analogue of PRGs, OWFs, and EFID [Gol90], respectively. [^1] EFI is a pair of efficiently generatable states that are statistically distinguishable but computationally not. EFID is its classical version: a pair of efficiently samplable classical distributions that are statistically distinguishable but computationally not. They seem to be weaker than OWFs [Kre21, KQST23, LMW23], yet can be used to construct several useful primitives such as private-key quantum money, SKE, commitments, multiparty computations, and digital signatures [MY22b, AQY22, BCQ23, Yan22].
@@
In this paper, we focus on OWSGs, which are a quantum analog of OWFs.222 There are a pure state output version [MY22b] and mixed state output version [MY22a] of OWSGs. We consider OWSGs with mixed state outputs as a default definition, but all of our positive results actually have pure state outputs. We adopt the mixed state version to make our negative results stronger. A OWSG is a quantum polynomial-time (QPT) algorithm that takes a bit string $k$ as input and outputs a quantum state $\phi_{k}$. The security is, roughly speaking, that no QPT adversary can find “preimage” of $\phi_{k}$. There are two differences from classical OWFs. First, because quantum states cannot be copied in general, we have to take care of the number of copies of $\phi_{k}$ that the adversary receives. In our definition, we assume that any polynomially many copies of $\phi_{k}$ are given to the adversary. Second, we have to explicitly define what do we mean by “preimage” unlike classical OWFs, because two different quantum states can have some non-zero overlap. We introduce a QPT verification algorithm that takes $k^{\prime}$ (which is the output of the adversary) and $\phi_{k}$ as input, and outputs $\top$ or $\bot$. [^3] If $\phi_{k}$ is a pure state, we can assume without loss of generality that the verification algorithm is the following one: project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ and output $\top$ if the projection is successful. Otherwise, output $\bot$.
@@
Recently, Cavalar, Goldin, Gray, Hall, Liu, and Pelecanos [${}^{+}$23], among many other results, raised the question of how short output length of OWSGs can be. They show that PRSGs with $\omega(\log\lambda)$-qubit outputs are also OWSGs, where $\lambda$ is the security parameter. [^4] A PRSG is a QPT algorithm that takes a bit string as input and outputs a quantum state whose polynomially-many copies are computationally indistinguishable from the same number of copies of Haar random states. Since it is known that OWFs imply PRSGs of arbitrary (QPT-computable) output length [BS20], their result gives OWSGs with $m(\lambda)$-qubit outputs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$ based on OWFs.555 We say that a function $m$ defined on $\mathbb{N}$ is QPT-computable if there is a QPT algorithm that computes $m(\lambda)$ on input $1^{\lambda}$ with probability $1$. Regarding a lower bound on the output length, they conjecture that there do not exist OWSGs with $O(\log\log\lambda)$-qubit outputs. However, they do not prove it and leave it open to study parameter regimes for which OWSGs do and do not exist.
@@ 1.1 Our Results
We prove that the conjecture of [${}^{+}$23] is true. Indeed, we prove a much stronger statement than their conjecture.
Theorem 1.1 ().
@@
Note that their conjecture is the impossibility of OWSGs with $O(\log\log\lambda)$-qubit outputs, but we prove the impossibility even for $O(\log\lambda)$-qubit outputs. The proof of the above theorem is quite simple. We simply consider a trivial attacker that just outputs a random input $k^{\prime}$ just ignoring the target state $\phi_{k}$. When $\phi_{k}$ is an $m$-qubit pure state, Welch bound [Wel74] (Lemma 2.4) directly implies that the above adversary has an advantage at least $2^{-m}$, which is inverse-polynomial when $m=O(\log\lambda)$. For the case of mixed state outputs, we need slightly more calculations, but we can still show that the above adversary has an inverse-polynomial advantage when $m=O(\log\lambda)$. The above theorem implies that the construction of $\omega(\log\lambda)$-output OWSGs based on OWFs from [BS20, ${}^{+}$23] is optimal in terms of the output length.
@@
The above theorem means that when the output length is $O(\log\lambda)$, there is an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. On the other hand, for OWFs with $O(\log\lambda)$-bit outputs, there is an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ for any polynomial ${\mathrm{poly}}$ (e.g., see [Gol01, Excercise 10 of Section 2]). Thus, it is natural to ask if such an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ exists for OWSGs with $O(\log\lambda)$-qubit outputs. First, we note that we cannot amplify the adversary’s advantage by repeating it many times in general. [^6] Note that such an amplification of the adversary’s advantage does not work even for OWFs. Indeed, if that was possible, then any weak OWF would be also a OWF. But we can easily construct a weak OWF that is not a OWF assuming the existence of OWF. Do not confuse this with amplification theorem that constructs OWFs from weak OWFs [Yao82]. The amplification theorem means that we can convert any weak OWF to a OWF, but it does not mean that any weak OWF is also a OWF. For example, suppose that an adversary given many copies of $\phi_{k}$ always outputs $k^{\prime}$ such that the verification algorithm accepts $(k^{\prime},\phi_{k})$ with probability $1/{\mathrm{poly}}(\lambda)$. Such an adversary has advantage $1/{\mathrm{poly}}(\lambda)$ against security of the OWSG, but one cannot gain anything by running it many times since it always outputs $k^{\prime}$ that has only $1/{\mathrm{poly}}(\lambda)$ acceptance probability. This leaves a possibility that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs if we weaken the security requirement. To better understand the tradeoffs between the output length and security, we introduce an $\epsilon$-OWSGs, which are a parameterized variant of OWSGs where the adversary’s advantage is only required to be at most $\epsilon$, and study parameter regimes for output length for which $\epsilon$-OWSGs do and do not exist.
@@
First, we study the regime of $\epsilon=\Theta(1)$. In this regime, we show that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs.
Theorem 1.2 ().
We show this theorem by using quantum fingerprinting [BCWdW01]. Roughly, quantum finger printing encodes an $\ell$-bit classical string $x$ into an $O(\log(\ell)\log(\eta^{-1}))$-qubit pure state |$\ket{h_{x}}$ in such a way that $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for any $x\neq x^{\prime}$. Our idea is to compose a OWF and quantum fingerprinting, i.e., our OWSG outputs |$\ket{h_{f(x)}}$ on input $x$ where $f$ is a OWF. When $\eta=\Theta(1)$, the length of the output is logarithmic in the output length of $f$. Thus, if we start from a OWF with ${\mathrm{polylog}}(\lambda)$-bit outputs, which exists assuming the existence of subexponentially secure OWFs, the output length of the above OWSG is $O(\log\log\lambda)$. We also show that the above construction preserves the one-wayness up to an additive constant error caused by quantum finger printing. This yields Theorem 1.2.
@@
We show the following lower bound for the output length in this regime.
Theorem 1.3 ().
@@
Our proof is inspired by the construction of quantum pseudorandom generators from short PRSGs [ALY23]. If their construction applies to the one-wayness, we can construct weak OWFs with $O(\log\lambda)$-bit output length, which is known to be impossible. Unfortunately, their result heavily relies on the pseudorandomness of outputs, and extending it to the OWSGs does not seem easy. Instead, we use the tomography and epsilon-nets [^7] An $\gamma$-net of $m$-qubit quantum states is a set of quantum states such that for any $m$-qubit quantum state, there exists an $\gamma$-close element in the set. We use $\gamma$ instead of $\epsilon$ to avoid confusion with the security of OWSGs. (Lemmata 2.6 and 2.7) to construct a randomized algorithm with classical outputs from the OWSGs, in which we can adapt the impossibility of weak OWFs with $O(\log\lambda)$-bit output length to break the original OWSGs. The output length is bounded by $O(\log\lambda)$ because the cardinality of $\gamma$-net in $m$-qubit space is bounded by $(C/\gamma)^{2^{2m}}$ for some constant $C$, which is polynomial for constant $\gamma$ and $m=\frac{1}{2}\log\log\lambda+O(1)$. For the OWSGs with pure state outputs, this can be improved to $\log\log\lambda+O(1).$
Next, we study the regime of $\epsilon=1-1/{\mathrm{poly}}(\lambda)$ for some polynomial ${\mathrm{poly}}$. We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs following [MY22a]. We show that slightly super-constant-qubit outputs are sufficient for weak OWSGs.
Theorem 1.4 ().
For any QPT-computable function $m(\lambda)=\omega(1)$, there exist weak OWSGs with $m(\lambda)$-qubit pure state outputs assuming the existence of exponentially secure injective OWFs that map $\lambda$-bit inputs to $O(\lambda)$-bit outputs.
The main observation for proving this theorem is the following. For an angle $\theta$, let |$\ket{+_{\theta}}:=\frac{1}{\sqrt{2}}\left(\ket{0}+e^{i\theta}\ket{1}\right)$. Then we can show that for any distinct $y,y^{\prime}\in[\lambda]$, we have $|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}=1-1/{\mathrm{poly}}(\lambda)$. This means that we can encode $\log\lambda$ classical bits into $1$ qubit if we allow the inner products to be as large as $1-1/{\mathrm{poly}}(\lambda)$. By repeating it, we can encode $t\log\lambda$ classical bits into $t$ qubits. By using this encoding instead of quantum fingerprinting in the construction for Theorem 1.2, we obtain weak OWSGs with $t$-qubit outputs starting from any injective OWFs with $t\log\lambda$-bit outputs. Assuming the existence of exponentially secure injective OWFs with linear expansion, there are injective OWFs with $t\log\lambda$-bit outputs for any super-constant $t$. This yields Theorem 1.4.
Finally, we show a matching lower bound for the output length of weak OWSGs.
Theorem 1.5 ().
@@
Our results are summarized in Table 1.
@@
(Assuming exp secure injective OWFs)

### v3

@@ Abstract
We study the output length of one-way state generators (OWSGs), their weaker variants, and EFIs.
@@
  • Inverse-polynomial-advantage OWSGs.
    Let $\epsilon$-OWSGs be a parameterized variant of OWSGs where a quantum polynomial-time adversary’s advantage is at most $\epsilon$. For any constant $c\in\mathbb{N}$, we construct $\lambda^{-c}$-OWSGs with $((c+1)\log\lambda+O(1))$-qubit outputs assuming the existence of OWFs. We show that this is almost tight by proving that there do not exist $\lambda^{-c}$-OWSGs with at most $(c\log\lambda-2)$-qubit outputs.
@@
    For any constant $\epsilon>0$, we construct $\epsilon$-OWSGs with $O(\log\log\lambda)$-qubit outputs assuming the existence of subexponentially secure OWFs. We show that this is almost tight by proving that there do not exist $O(1)$-OWSGs with $((\log\log\lambda)/2+O(1))$-qubit outputs.
@@
    We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs. We construct weak OWSGs with $m$-qubit outputs for any $m=\omega(1)$ assuming the existence of exponentially secure OWFs with linear expansion. We show that this is tight by proving that there do not exist weak OWSGs with $O(1)$-qubit outputs.
  • EFIs.
    We show that there do not exist $O(\log\lambda)$-qubit EFIs. We show that this is tight by proving that there exist $m(\lambda)$-qubit EFIs for any $m(\lambda)=\omega(\log\lambda)$ assuming the existence of exponentially secure pseudorandom generators.
@@ 1 Introduction
In classical cryptography, the existence of one-way functions (OWFs) is the minimum assumption, because many primitives (such as secret-key encryption (SKE), commitments, zero-knowledge, pseudorandom generators (PRGs), pseudorandom functions (PRFs), and digital signatures) are equivalent to OWFs, and almost all cryptographic primitives imply OWFs [LR86, IL89, ILL89]. On the other hand, in quantum cryptography where quantum computing and quantum communications are available, recent studies suggest that OWFs would not be the minimum assumption, and many new replacements of OWFs have been introduced. For example, pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22b, MY22a], and EFIs [BCQ23] are quantum analogue of PRGs, OWFs, and EFIDs [Gol90], respectively. They seem to be weaker than OWFs [Kre21, KQST23, LMW23], yet can be used to construct several useful primitives such as private-key quantum money, SKE, commitments, multiparty computations, and digital signatures [MY22b, AQY22, BCQ23, Yan22].
@@
OWSGs are a quantum analog of OWFs.111 There are a pure state output version [MY22b] and mixed state output version [MY22a] of OWSGs. We consider OWSGs with mixed state outputs as a default definition, but all of our positive results actually have pure state outputs. We adopt the mixed state version to make our negative results stronger. A OWSG is a quantum polynomial-time (QPT) algorithm that takes a bit string $k$ as input and outputs a quantum state $\phi_{k}$. The security is, roughly speaking, that no QPT adversary can find “preimage” of $\phi_{k}$. There are two differences from classical OWFs. First, because quantum states cannot be copied in general, we have to take care of the number of copies of $\phi_{k}$ that the adversary receives. In our definition, we assume that any polynomially many copies of $\phi_{k}$ are given to the adversary. Second, we have to explicitly define what do we mean by “preimage” unlike classical OWFs, because two different quantum states can have some non-zero overlap. We introduce a QPT verification algorithm that takes $k^{\prime}$ (which is the output of the adversary) and $\phi_{k}$ as input, and outputs $\top$ or $\bot$. [^2] If $\phi_{k}$ is a pure state, we can assume without loss of generality that the verification algorithm is the following one: project |$\ket{\phi_{k}}$ onto |$\ket{\phi_{k^{\prime}}}$ and output $\top$ if the projection is successful. Otherwise, output $\bot$.
@@
Recently, Cavalar, Goldin, Gray, Hall, Liu, and Pelecanos [CGG+23], among many other results, raised the question of how short output length of OWSGs can be. They show that PRSGs with $\omega(\log\lambda)$-qubit outputs are also OWSGs, where $\lambda$ is the security parameter. [^3] A PRSG is a QPT algorithm that takes a bit string as input and outputs a quantum state whose polynomially-many copies are computationally indistinguishable from the same number of copies of Haar random states. Since it is known that OWFs imply PRSGs of arbitrary (QPT-computable) output length [BS20], their result gives OWSGs with $m(\lambda)$-qubit outputs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$ based on OWFs.444 We say that a function $m$ defined on $\mathbb{N}$ is QPT-computable if there is a QPT algorithm that computes $m(\lambda)$ on input $1^{\lambda}$ with probability $1$. Regarding a lower bound on the output length, they conjecture that there do not exist OWSGs with $O(\log\log\lambda)$-qubit outputs. However, they do not prove it and leave it open to study parameter regimes for which OWSGs do and do not exist.
An EFI is a pair of efficiently generatable states that are statistically far but computationally indistinguishable. While relationships between EFIs and other primitives have been actively studied [MY22a, BCQ23], to our knowledge, there is no existing work that studies how many qubits are needed for EFIs.
@@ 1.1 Our Results
Lower Bounds for ($\epsilon$-)OWSGs.
We prove that the conjecture of [CGG+23] about the lower bound for the output length of OWSGs is true. Indeed, we prove a much stronger statement than their conjecture.
Theorem 1.1.
@@
Note that their conjecture is the impossibility of OWSGs with $O(\log\log\lambda)$-qubit outputs, but we prove the impossibility even for $O(\log\lambda)$-qubit outputs. The proof of the above theorem is quite simple. We simply consider a trivial attacker that just outputs a random input $k^{\prime}$ just ignoring the target state $\phi_{k}$. When $\phi_{k}$ is an $m$-qubit pure state, Welch bound [Wel74] (Lemma 2.8) directly implies that the above adversary has an advantage at least $2^{-m}$, which is inverse-polynomial when $m=O(\log\lambda)$. For the case of mixed state outputs, we need slightly more calculations, but we can still show that the above adversary has an inverse-polynomial advantage when $m=O(\log\lambda)$.
@@
The above theorem means that when the output length is $O(\log\lambda)$, there is an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. On the other hand, for OWFs with $O(\log\lambda)$-bit outputs, there is an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ for any polynomial ${\mathrm{poly}}$ (e.g., see [Gol01, Excercise 10 of Section 2]). Thus, it is natural to ask if such an adversary with advantage $1-1/{\mathrm{poly}}(\lambda)$ exists for OWSGs with $O(\log\lambda)$-qubit outputs. First, we note that we cannot amplify the adversary’s advantage by repeating it many times in general. [^5] Note that such an amplification of the adversary’s advantage does not work even for OWFs. Indeed, if that was possible, then any weak OWF would be also a OWF. But we can easily construct a weak OWF that is not a OWF assuming the existence of OWF. Do not confuse this with the amplification theorem that constructs OWFs from weak OWFs [Yao82]. The amplification theorem means that we can convert any weak OWF to a OWF, but it does not mean that any weak OWF is also a OWF. For example, suppose that an adversary is given many copies of $\phi_{k}$ always outputs $k^{\prime}$ such that the verification algorithm accepts $(k^{\prime},\phi_{k})$ with probability $1/{\mathrm{poly}}(\lambda)$. Such an adversary has an advantage $1/{\mathrm{poly}}(\lambda)$ against the security of the OWSG, but one cannot gain anything by running it many times since it always outputs $k^{\prime}$ that has only $1/{\mathrm{poly}}(\lambda)$ acceptance probability. This leaves a possibility that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs if we weaken the security requirement. To better understand the tradeoffs between the output length and security, we introduce an $\epsilon$-OWSGs, which are a parameterized variant of OWSGs where the adversary’s advantage is only required to be at most $\epsilon$, and study parameter regimes for output length for which $\epsilon$-OWSGs do and do not exist.
@@
First, we study the regime of $\epsilon=\lambda^{-\Theta(1)}$. We show that the proof of Theorem 1.1 naturally extends to this setting to obtain the following theorem.
Theorem 1.2.
For any constant $c\in\mathbb{N}$, there do not exist $\lambda^{-c}$-OWSGs of which output length is at most $c\log\lambda-2$.
@@
Next, we study the regime of $\epsilon=O(1)$. We show the following lower bound.
Theorem 1.3.
@@
In this regime, the idea for the proofs of Theorems 1.1 and 1.2 does not work since Welch bound only gives an adversary with advantage $1/{\mathrm{poly}}(\lambda)$. Thus, we rely on a different idea. Our proof is inspired by the construction of quantum pseudorandom generators from short PRSGs [ALY23]. If their construction applies to the one-wayness, we can construct weak OWFs with $O(\log\lambda)$-bit output length, which is known to be impossible. Unfortunately, their result heavily relies on the pseudorandomness of outputs, and extending it to the OWSGs does not seem easy. Instead, we use the tomography and epsilon-nets [^6] An $\gamma$-net of $m$-qubit quantum states is a set of quantum states such that for any $m$-qubit quantum state, there exists an $\gamma$-close element in the set. We use $\gamma$ instead of $\epsilon$ to avoid confusion with the security of OWSGs. (Lemmata 2.10 and 2.12) to construct a randomized algorithm with classical outputs from the OWSGs, in which we can adapt the impossibility of weak OWFs with $O(\log\lambda)$-bit output length to break the original OWSGs. The output length is bounded by $O(\log\lambda)$ because the cardinality of $\gamma$-net in $m$-qubit space is bounded by $(C/\gamma)^{2^{2m}}$ for some constant $C$, which is polynomial for constant $\gamma$ and $m=\frac{1}{2}\log\log\lambda+O(1)$. For the OWSGs with pure state outputs, this can be improved to $\log\log\lambda+O(1).$
Finally, we study the regime of $\epsilon=1-1/{\mathrm{poly}}(\lambda)$ for some polynomial ${\mathrm{poly}}$. We refer to $(1-1/{\mathrm{poly}}(\lambda))$-OWSGs as weak OWSGs following [MY22a]. We show the following lower bound.
Theorem 1.4.
@@
Upper Bounds for ($\epsilon$)-OWSGs.
We show the above lower bounds are almost tight by showing almost matching upper bounds.
As already mentioned, the combination of existing works [BS20, CGG+23] gives $m(\lambda)$-qubit output OWSGs based on OWFs for any QPT-computable function $m(\lambda)=\omega(\log\lambda)$. This means that Theorem 3.1 is tight. We show that their construction can be extended to $\lambda^{-\Theta(1)}$-OWSGs to obtain the following theorem.
Theorem 1.5.
For any constant $c\in\mathbb{N}$, there exist $\lambda^{-c}$-OWSGs with $((c+1)\log\lambda+O(1))$-qubit pure state outputs assuming the existence of OWFs.
Comparing this with Theorem 1.2, we can see that it is almost tight though there is a small gap between $c\log\lambda-2$ and $(c+1)\log\lambda+O(1)$. The above theorem means that we can circumvent the barrier of $O(\log\lambda)$-qubit outputs for OWSGs (Theorem 1.1) if we slightly weaken (${\mathsf{negl}}(\lambda)$-)OWSGs to $\lambda^{-\Theta(1)}$-OWSGs.
Next, we show an upper bound for the output length of $\Theta(1)$-OWSGs.
Theorem 1.6.
Comparing this with Theorem 1.3, we can see that it is tight ignoring the constant factor for $\log\log\lambda$. We show this theorem by using quantum fingerprinting [BCWdW01]. Roughly, quantum finger printing encodes an $\ell$-bit classical string $x$ into an $O(\log(\ell)\log(\eta^{-1}))$-qubit pure state |$\ket{h_{x}}$ in such a way that $|\langle h_{x}|h_{x^{\prime}}\rangle|\leq\eta$ for any $x\neq x^{\prime}$. Our idea is to compose a OWF and quantum fingerprinting, i.e., our OWSG outputs |$\ket{h_{f(x)}}$ on input $x$ where $f$ is a OWF. When $\eta=\Theta(1)$, the length of the output is logarithmic in the output length of $f$. Thus, if we start from a OWF with ${\mathrm{polylog}}(\lambda)$-bit outputs, which exists assuming the existence of subexponentially secure OWFs, the output length of the above OWSG is $O(\log\log\lambda)$. We also show that the above construction preserves the one-wayness up to an additive constant error caused by quantum finger printing. This yields Theorem 1.6.
Finally, we show an upper bound for the output length of weak OWSGs.
Theorem 1.7.
For any QPT-computable function $m(\lambda)=\omega(1)$, there exist weak OWSGs with $m(\lambda)$-qubit pure state outputs assuming the existence of exponentially secure OWFs that map $\lambda$-bit inputs to $O(\lambda)$-bit outputs.
Comparing this with Theorem 1.4, we can see that this is tight. The main observation for proving this theorem is the following. For an angle $\theta$, let |$\ket{+_{\theta}}:=\frac{1}{\sqrt{2}}\left(\ket{0}+e^{i\theta}\ket{1}\right)$. Then we can show that for any distinct $y,y^{\prime}\in[\lambda]$, we have $|\langle+_{2\pi y/\lambda}|+_{2\pi y^{\prime}/\lambda}\rangle|\leq\sqrt{\frac{1+\cos(2\pi/\lambda)}{2}}=1-1/{\mathrm{poly}}(\lambda)$. This means that we can encode $\log\lambda$ classical bits into $1$ qubit if we allow the inner products to be as large as $1-1/{\mathrm{poly}}(\lambda)$. By repeating it, we can encode $t\log\lambda$ classical bits into $t$ qubits. By using this encoding instead of quantum fingerprinting in the construction for Theorem 1.6, we obtain weak OWSGs with $t$-qubit outputs starting from any OWFs with $t\log\lambda$-bit outputs. Assuming the existence of exponentially secure OWFs with linear expansion, there are OWFs with $t\log\lambda$-bit outputs for any super-constant $t$. This yields Theorem 1.7.
Our results on upper and lower bounds for ($\epsilon$-)OWSGs are summarized in Table 1.
@@
$\lambda^{-c}$ $(c+1)\log\lambda+O(1)$ [Theorem 1.5] $c\log\lambda-2$ [Theorem 1.2]
$(c=\Theta(1))$ (Assuming OWFs)
@@
(Assuming exp secure OWFs)
EFIs.
We show matching lower and upper bounds for the length of EFIs. The lower bound is stated as follows.
Theorem 1.8.
There do not exist $O(\log\lambda)$-qubit EFIs.
We remark that the proof of this theorem is straightforward in the non-uniform setting: Since the two states that form an EFI are statistically far, there exists a measurement that distinguishes them with advantage $1/{\mathrm{poly}}(\lambda)$. If they are $O(\log\lambda)$-qubit states, the optimal distinguishing measurement can be described as a ${\mathrm{poly}}(\lambda)$-bit classical string. Thus, a non-uniform adversary that takes the description of the measurement as advice can break the computational indistinguishability. To extend it to the uniform setting, our idea is to use tomography. [^7] We thank Fermi Ma for suggesting the use of tomography. Roughly, we show that a uniform adversary can use tomography on the two states to efficiently approximate the optimal distinguishing measurement.
We show a matching upper bound as follows.
Theorem 1.9.
If exponentially secure PRGs exist, then $m(\lambda)$-qubit EFIs exist for any QPT computable function $m(\lambda)=\omega(\log\lambda)$.
The proof is almost obvious. By using an exponentially secure PRG $G:\{0,1\}^{m/2}\rightarrow\{0,1\}^{m}$, let $\rho_{0}:=\frac{1}{2^{m/2}}\sum_{x\in\{0,1\}^{m/2}}\ket{G(x)}\bra{G(x)}$ | and $\rho_{1}:=\frac{1}{2^{m}}\sum_{y\in\{0,1\}^{m}}\ket{y}\bra{y}$ |. Then $\rho_{0}$ and $\rho_{1}$ are computationally indistinguishable by exponential security of the PRG and $m(\lambda)=\omega(\log\lambda)$, but they are statistically far since the image size of $G$ is at most $2^{m/2}\ll 2^{m}$.
Quantum bit commitments.
It is known that EFIs exist if and only if quantum bit commitments exist [Yan22, BCQ23]. However, the conversion between them does not preserve the number of qubits, and thus the lower and upper bounds for the length of EFIs do not necessarily imply similar bounds for quantum bit commitments. To complement it, we also show similar lower and upper bounds for the commitment length (i.e., the number of qubits of the commitment register) of quantum bit commitments. That is, we show a $O(\log\lambda)$ lower bound and $\omega(\log\lambda)$ upper bound assuming the existence of exponentially secure PRGs. See Appendix A for the details.